OpenStudy (anonymous):

i need help with this question what is 2 1/5 + 1 1/10? a. 3 1/5 b. 3 3/5 c. 3 1/10 d. 3 3/10

OpenStudy (kamibug):

2 + 1 = 3 2/10 + 1/10 = 3/10 3 3/10

OpenStudy (anonymous):

thanks